A lead bullet of mass 50gram moving witha velocity of 10m per Sec strikes a target. If 40%is retained by the bullet. Caclulate the raise in it`s temperature where Specific heat equal to 4.2 joule per cal.

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Energy = 1/2 * 0.05 * 10² = 2.5 joule
Now
m* s * delta T = 2.5
0.05 * 4.2 * delta T = 2.5
delta T = 2.5/4.2* 0.05 = 11.904
Hence its temperature increases by 11.9°C.
 
Regards
Arun (askIITians forum expert)